OpenStudy (anonymous):

A bag has 1 red marble, 4 blue marbles, and 3 green marbles. Peter draws a marble randomly from the bag, replaces it, and then draws another marble randomly. What is the probability of drawing 2 blue marbles in a row? Explain your answer.

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

oky so i guess \[\rm \frac{ blue~marbles }{ total ~number~of marbles }\]

Nnesha (nnesha):

ahhh i'm not good 't probability stuff ?/>

Nnesha (nnesha):

@jim_thompson5910

OpenStudy (anonymous):

ITS OK THANKS

jimthompson5910 (jim_thompson5910):

@Nnesha you're on the right track

OpenStudy (anonymous):

Would it be 2/8

jimthompson5910 (jim_thompson5910):

you need to figure out how many blue marbles there are, out of how many total probability of picking blue = (# of blue)/(# total) since replacements are made, this means that the probability does not change on the second draw

jimthompson5910 (jim_thompson5910):

probability of picking 2 blue = (probability of picking 1 blue)*(probability of picking 1 blue)

OpenStudy (anonymous):

so im kinda lost...so it would be 2= 1x1?

jimthompson5910 (jim_thompson5910):

we have 4 blue out of 8 total so you agree that the probability of picking one blue is 4/8 = 1/2 right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

the probability of picking 2 blue is equal to P(2 blue) = P(1 blue AND 1 blue) P(2 blue) = P(1 blue)*P(1 blue) P(2 blue) = (1/2)*(1/2) P(2 blue) = (1*1)/(2*2) P(2 blue) = 1/4

jimthompson5910 (jim_thompson5910):

this rule only works because replacements are made and each drawing is independent

OpenStudy (anonymous):

OK so what should i write down? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Basically those steps and the answer. But in your own words and format that's unique to you. Hopefully you see how I got that answer?

OpenStudy (anonymous):

yes thank u!

jimthompson5910 (jim_thompson5910):

you're welcome

Nnesha (nnesha):

thanks @jim_thompson5910 o^_^o